OpenStudy (anonymous):

Find the area of a trapezoid with a height of 4 m and bases of 12 m and 19 m.

OpenStudy (anonymous):

A=(h/2)(B1+B2) just plug in

OpenStudy (anonymous):

like this A = 1/2h(b1 + b2) Now multiply both sides by 2 to remove the fraction on the right: 2A = h(b1 + b2) Now divide both sides by h: 2A/h = b1 + b2 Now substract b2 from both sides: 2A/h - b2 = b1

OpenStudy (anonymous):

is that good

OpenStudy (anonymous):

can someone tell if i did this good

OpenStudy (anonymous):

sorry,was busy: u want 2 find A so A=(4/2)(12+19)=2(31)=62 m^2

OpenStudy (anonymous):

thank i did it